I need help solving this for my 9th grade son
Find the line of best fit between the following points:
(1,5) and (3,8)
a) 1.5
b) 3.5
(his answers)
line: ____?_____
His answer 1.5 and 3.5 came from calculator instructions, he needs the step by step instructions for solving (without a calculator).
Any help is greatly appreciated.

Answer by scott8148(6628)

the a) answer is the slope and the b) answer is the y-intercept

find the slope first
___ the difference in the y-coordinates, divided by, the difference in the x-coordinates (rise over run)
___ make sure you take the difference in the same direction for both coordinates

in this case the slope (m) equals ___ (8 - 5) / (3 - 1) = 3/2 = 1.5

using the slope-intercept form of the line (y = mx + b), substitute one of the points to find the y-intercept (b)

5 = (1.5)1 + b ___ 3.5 = b

so the line (in slope-intercept form) is ___ y = (1.5)x + 3.5

yes, the calculator does work